A motorcycle manufacturer produces three different models: the Avalon, the Durango, and the Roadtripper. Production restrictions require it to make, on a monthly basis, 10 more Roadtrippers than the total of the other two models, and twice as many Durangos as Avalons. The shop must produce a total of 490 cycles per month. How many cycles of each type should be made per month?
I just don't know how to set it up!!!! Thanks in advance!

Answer by checkley71(8403)

A+D+R=490 GIVEN
R=10+(A+D) GIVEN
D=2A GIVEN
THEN R=10(A+2A) OR
R=10+(A+2A)
R=10+3A
THEN A+(2A)+(10+3A)=490
A+2A+10+3A=490
6A=490-10
6A=480
A=480/6
A=80 AVALONS
D=2*80
D=160 DURANGOS
R=10+(80+160)
R=10+240
PROOF
R=250
A+D+R=490
80+160+250=490
490=490